### v1

# WHALE-FL: Wireless and Heterogeneity Aware Latency Efficient Federated Learning over Mobile Devices via Adaptive Subnetwork Scheduling

Abstract
As a popular distributed learning paradigm, federated learning (FL) over mobile devices fosters numerous applications, while their practical deployment is hindered by participating devices’ computing and communication heterogeneity. Some pioneering research efforts proposed to extract subnetworks from the global model, and assign as large a subnetwork as possible to the device for local training based on its full computing and communications capacity. Although such fixed size subnetwork assignment enables FL training over heterogeneous mobile devices, it is unaware of (i) the dynamic changes of devices’ communication and computing conditions and (ii) FL training progress and its dynamic requirements of local training contributions, both of which may cause very long FL training delay. Motivated by those dynamics, in this paper, we develop a wireless and heterogeneity aware latency efficient FL (WHALE-FL) approach to accelerate FL training through adaptive subnetwork scheduling. Instead of sticking to the fixed size subnetwork, WHALE-FL introduces a novel subnetwork selection utility function to capture device and FL training dynamics, and guides the mobile device to adaptively select the subnetwork size for local training based on (a) its computing and communication capacity, (b) its dynamic computing and/or communication conditions, and (c) FL training status and its corresponding requirements for local training contributions. Our evaluation shows that, compared with peer designs, WHALE-FL effectively accelerates FL training without sacrificing learning accuracy.
[…]
1 Introduction
Federated Learning (FL) [20] recently experienced a notable evolution, expanding its scope from conventional data center environments to harness the potential of mobile devices [15, 18, 4]. This shift has been propelled by the continuous advancements in hardware, empowering mobile devices like the NVIDIA Xavier, Galaxy Note20, iPad Pro, MacBook laptops, etc. with increasingly robust on-device computing capabilities tailored for local training. With the collective intelligence of ending edge devices and FL’s fundamental principle of preserving data privacy, FL over mobile devices has paved the way for a diverse spectrum of innovative mobile applications, including keyboard predictions [11], smart home hazard detection [24], health event detection [3], and so on.
While FL over mobile device has great potentials, its practical deployment faces significant challenges due to the inherent heterogeneity among real-world mobile devices, varying in computing capability, wireless conditions and local data distribution [14]. Existing FL studies often assume the model-homogeneous setting, where global and local models share identical architectures across all clients. However, as devices are forced to train models within their individual capability, developers have to choose between excluding low-tier devices, introducing training bias [2], or maintaining a low-complexity global model to accommodate all clients, resulting in degraded accuracy [6, 23]. The trend towards large models like Transformers [19] exacerbates the issue, hindering their training on mobile devices. Furthermore, unlike GPU clusters with stable high-speed Internet connections, mobile devices’ computing resources are constrained and heterogeneous and their wireless transmissions are relatively slow and dynamic, both of which lead to huge latency in FL training [5] and may severely degrade the performance of associated applications.
To address the limitations of model-homogeneous FL, researchers have recently studied how to enable heterogeneous models across the server and mobile clients during FL training. Subnetwork training, exemplified by pioneering approaches like Federated Dropout [22], HeteroFL [9], and FjORD [13], exhibits efficacy by having mobile devices train smaller subnetworks extracted from the large global server model. These designs offer solutions to aggregating diverse devices’ subnetworks. By tailoring subnetwork architecture for the individual device, subnetwork training can ensure compatibility with mobile devices owning heterogeneous computing and communications capability. However, a prevalent challenge in current subnetwork approaches lies in their static fixed-size subnetwork assignment policy. Such a policy may fail to realize the full potential of subnetwork based training, mainly due to the unawareness of system dynamics (i.e., computing and communications dynamics) and FL training dynamics.
System dynamics encompass the time-varying computing loads of devices’ background applications and the fluctuating wireless communication conditions across FL training rounds, which affects the sizes of subnetworks that a mobile device can support over rounds. Since most modern mobile devices (e.g., smartphones) participating in FL training have the ability to run multiple tasks (e.g., video streaming, image processing, and social media updates [1]) simultaneously, the dynamic orchestration of CPU/GPU resources across these concurrent activities results in the fluctuations in computing power and available memory for FL tasks, consequently impacting the supported subnetwork sizes for on-device computing. Similarly, wireless communications dynamics caused by users’ mobility, wireless channel fading, etc. lead to dynamic transmission rates, which directly affect candidate subneworks sizes that a mobile device can support for local model updates.
FL training dynamics represents FL convergence’s dynamic requirements for the contributions from local training at different training stages, which implicitly affects participating devices’ selections on subnetwork sizes. As FL training starts from scratches, any contributions from local training is helpful at the early training stage. Then, when FL training steps into the middle stage, it requires more accurate local training results to converge the global model. Finally, at the late stage, i.e., when FL training is close to the convergence, most participating devices have already made sufficient contributions to the global model and their local training may only have marginal benefits for global model’s convergence.
We observe that failing to capture system or FL training dynamics and always using the possible largest-sized subnetworks under devices’ full capabilities may significantly prolong FL training process. Different from prior static fixed-size subnetwork assignment methods, in this paper, we propose a wireless and heterogeneity aware latency efficient FL (WHALE-FL) approach to accelerate FL training via adaptive subnetwork scheduling. WHALE-FL characterizes system dynamics and FL training dynamics and tailors appropriate sized subnetworks for heterogeneous mobile devices under dynamic computing/wireless environments at different FL training stages. As far as we know, WHALE-FL is the first paper that converts static fixed-size subnetwork allocation (e.g., HeteroFL [9], FjORD [13], etc.) into dynamic/adaptive subnetwork scheduling for each device by jointly considering system heterogeneity and FL training dynamics, and conducts system level experiments for validation. Our salient contributions are summarized as follows.
  • We design a novel subnetwork selection utility function to capture system and FL training dynamics, guiding mobile devices to adaptively size their subnetworks for local training based on the time-varying computing/communication capacity and FL training status.
  • We develop a WHALE-FL prototype and evaluate its performance with extensive experiments. The experimental results validate that WHALE-FL can remarkably reduce the latency for FL training over heterogeneous mobile devices without sacrificing learning accuracy.
2 Preliminary
2.1 FL over Heterogeneous Mobile Devices
Consider that $M$ mobile devices in a wireless network collaboratively engage in FL to train a deep neural network on locally distributed data sets $\{D_{1},\cdots,D_{i},\cdots,D_{M}\}$. Their local models are parameterized by $\{W_{1},\cdots,W_{i},\cdots,W_{M}\}$, which are updated using stochastic gradient descents [21] on the local data samples through local training. The server collects the local model updates and aggregates them into a global model $W_{g}$ using model averaging [20, 17]. This aggregation occurs over multiple communication rounds, with the global model at the $r$-th round denoted as $W_{g}^{r}=\frac{1}{M}\sum_{m=1}^{M}W_{m}^{r}$. In the subsequent training round, $W_{g}^{r}$ is transmitted to mobile devices, and their local models are updated as $W_{i}^{r+1}=W_{g}^{r}$. This process repeats until FL converges, while system heterogeneity (communications and computing) among mobile devices incurs huge training latency and significantly slows down FL convergence.
2.2 FL with Subnetwork Extraction
To address system heterogeneity issue in FL training, subnetwork method was introduced in [9], which extracts different sized subnetworks from global model.
Let $\mathcal{W}^{P}=\{W^{1},W^{2},\cdots,W^{p},\cdots,W^{P}\}$ be a collection of candidate subnetworks to be selected by mobile devices for local training, where $P$ complexity/size levels are considered. A lower size level $p$ corresponds to a larger-sized subnetwork, and $W^{P}$ is the smallest subnetwork for selection, i.e., $W^{P}\subset W^{P-1}\subset\cdots\subset W^{1}$. We follow the same approach as illustrated in [9] to extract subnetworks from the global model by shrinking the width of hidden channel with specific ratios. Let $s\in(0,1]$ be the hidden channel shrinkage ratio. Then, we have $|W^{p}|/|W_{g}|=|W^{p}|/|W^{1}|=s^{2(p-1)}$. With this construction, different sized subnetworks can be assigned to participating mobile devices according to their corresponding capabilities. Suppose that the number of devices in each subnetwork size level is $\{M_{1},\cdots,M_{P}\}$. The server has to aggregate the heterogeneous subnetworks in every training round. As demonstrated in [9], the global aggregation is conducted as follows.
$W_{g}=W_{g}^{1}$ $=W_{g}^{P}\cup(W_{g}^{P-1}\backslash W_{g}^{P})\cup\cdots\cup(W_{g}^{1}\backslash W_{g}^{2})$
$=W_{g}^{P}\cup\bigcup^{P}_{p=2}W_{g}^{p-1}\setminus W_{g}^{p},$ (1)
where
$W_{g}^{P}=\frac{1}{M}$ $\sum\limits_{m=1}^{M}W_{m}^{P},$
$W_{g}^{p-1}\backslash W_{g}^{p}$ $=\frac{1}{M-M_{p:P}}\sum\limits_{m=1}^{M-M_{p:P}}W_{m}^{p-1}\backslash W_{m}^{p},\forall p\in[2,P].$
In this way, each parameter is averaged from those devices whose assigned subnetwork contains that specific parameter, which enables the global aggregation and FL training with different sizes of subnetworks. Although the subnetwork method in [9] alleviates system heterogeneity issue, it is a fixed policy. It cannot capture the dynamic changes of wireless transmission/on-device computing conditions, or the dynamic requirements of contributions from local training at different FL training stages, either of which may result in a huge FL training latency.
3 Motivation
Unawareness of system dynamics. Traditional subnetwork assignment (e.g., HeteroFL [9]) is fixed, which is based on the participating mobile device’s maximum system capability (i.e., computing + communications), while ignoring the dynamic changes of the device’s computing and communication conditions. Such an unawareness may lead to poor subnetwork assignment decisions and significantly delay the FL training process. For instance, a mobile device capable of computing a full-sized model may be experiencing a bad wireless access (e.g., 4G/LTE) or running some computing intensive background applications (e.g., GPU intensive gaming) in a certain training round. In this case, the fixed full-sized subnetwork assignment may become this device a straggler and cause a big latency in FL training. Thus, an adaptive subnetwork scheduling aware of system (computing + communication) dynamics is in need.
Unawareness of FL training dynamics. The fixed subnetwork assignment is unaware of FL training progress and its dynamic requirements of learning contributions from local mobile devices. Specifically, in early FL training stage, FL training starts from scratches and any contributions from any device’s local training will be helpful, regardless of the subnetwork size. At this stage, using small-sized subnetworks can expedite on-device computing and wireless transmissions of local model updates, which helps to reduce the FL training latency. Then, as FL training proceeds, say the middle stage, more and more accurate local model updates are needed for global training model to converge. Thus, only using small-sized subnetwork for speed while sacrificing local learning performance is not a wise option for mobile devices, since it may slow down or even fail FL convergence. Further, when FL training is close to the convergence (i.e., the late stage), most mobile devices have already made substantial/enough contributions to the global model. As for those devices, sticking to the large/full-sized subnetwork for local training has very limited learning benefits for FL convergence, while some computing/communications constrained ones among those devices may incur huge training latency or even become stragglers. Therefore, it is necessary to develop an adaptive subnetwork scheduling that captures FL training dynamics, recognizes computing/communication constraints and selects right sized subnetworks for local training, to improve the delay efficiency in FL training over mobile devices.
4 WHALE-FL Design
Aiming to reduce FL training latency, WHALE-FL entitles mobile devices to distributedly schedule different sizes of subnetworks for local training, adapting to their system dynamics and FL training dynamics. To capture those dynamics, WHALE-FL presents a novel adaptive subnetwork selection utility function jointly considering system efficiency and FL training efficiency. Moreover, WHALE-FL provides a normalization procedure to convert the calculated subnetwork selection utility values to discrete size levels of subnetowrks for mobile devices’ local scheduling decisions.
Figure 1: The conceptual sketch of WHALE-FL.
4.1 Adaptive Subnetwork Selection Utility
WHALE-FL’s adaptive subnetwork selection performance hinges on two critical aspects: system efficiency and training efficiency. System efficiency encompasses the duration of each training round, including local computing and model updates time consumption. Training efficiency gauges the local training’s contributions to global convergence. The fluctuating wireless conditions and available computing resources of devices, as well as their training progress with local data, collectively determine the system and training efficiency, forming what we term as adaptive subnetwork selection utility.
To accelerate FL training without sacrificing learning accuracy, it is critical to trade-off system and training efficiencies to select the appropriate subnetwork size for individual device’s local training per round. Briefly, WHALE-FL favors system efficiency over training efficiency at the early stage of FL training, and tends to schedule small-sized subnetworks for devices’ local training. While FL training steps into the middle stage, if more accurate local training is needed for FL convergence, WHALE-FL prefers training efficiency to system efficiency and schedules to adaptively increase the size of subnetworks for participating mobile devices. Otherwise, WHALE-FL prioritizes system efficiency over training efficiency. When FL is close to convergence, WHALE-FL jointly considers system and training efficiencies, and gradually decreases the size of subnetworks for local training, given the fact that most devices have contributed enough to global model and it is unnecessary to keep large-sized subnetworks for local training.
System efficiency utility. We define the system efficiency ($SE_{i,r}$) for any given client $i$ in the $r$-th round based on its wireless transmission rate and available computing resources at that time, which is calculated as follows:
$SE_{i,r}=\frac{1}{T_{i,r}^{tr}+T_{i,r}^{co}},$ (3)
where $T_{i,r}^{tr}$ and $T_{i,r}^{co}$ are the transmission delay and the computing delay, respectively, for the unit/smallest subnetwork. We assume that the wireless transmission rates and available computing resources dynamically change over rounds, but are relatively stable within a FL training round. Thus, given a learning task, transmission and computing workloads for the unit subnetwork are fixed, and $T_{i,r}^{tr}$ and $T_{i,r}^{co}$ can be easily estimated for device $i$ in the $r$-th round. A higher $SE_{i,r}$ enables devices to opt for larger subnetwork sizes for local training within this round, and vice versa. The formulation in Eqn. (3) comprehensively covers the system efficiency for communication delay dominant cases (i.e., slow transmissions & fast computing), computing delay dominant cases (i.e., fast transmissions & slow computing), and communication-computing comparable cases.
Training efficiency utility. WHALE-FL employs training efficiency utility to characterize FL training dynamics. WHALE-FL leverages training loss that measures the estimation error between model predictions and the ground truth to identify the current training status. Let $l_{i,r}$ denote the average training loss on the samples over all the local iterations in the $r$-th round, which can be automatically generated during on-device training with negligible collection overhead. In an adaptive update manner, we define the training efficiency $TE_{i,r}$ for device $i$ in the $r$-th round as follows:
$TE_{i,r}=\quad TE_{i,r-1}\quad+$
$\gamma\cdot[2\cdot\mathbbm{1}(l_{i,r-1}\geq\alpha\cdot l_{target})-1]\cdot\mathbbm{1}(\Delta l_{i,r-1}\leq\Delta_{th}),$ (4)
where $l_{target}$ denotes the target loss depending on the specific FL training task, $\mathbbm{1}(x)$ is an indicator function that takes value $1$ if the statement $x$ is true and 00 otherwise, $\alpha\geq 1$ is a developer-defined constant governing the delineation of training stages, and $\gamma\geq 0$ is a hyper-parameter controlling the step-size of utility increase or decrease compared to the previous round. Besides, $\Delta l_{i,r}=l_{i,r}-l_{i,r-1}$ denotes the gradient of local training loss at device $i$ in the $r$-th round, and $\Delta_{th}$ is the developer-specified threshold for local training loss gradient. Here, $\mathbbm{1}(l_{i,r-1}\geq\alpha\cdot l_{target})$ represents whether the FL training is close to the convergence, and $\mathbbm{1}(\Delta l_{i,r-1}\leq\Delta_{th})$ represents whether device $i$’s local training can make good statistical contributions [14] to FL training.
At the early FL training stage, $(l_{i,r-1}\geq\alpha\cdot l_{target})$ and $(\Delta l_{i,r-1}>\Delta_{th})$, which indicates that device $i$ makes sufficient contributions to FL training, and should keep the subnetwork size from training efficiency perspective. Thus, following Eqn. (4.1), $TE_{i,r}=TE_{i,r-1}$. At the middle FL training stage, there are two cases: Case I: $(l_{i,r-1}\geq\alpha\cdot l_{target})$ and $(\Delta l_{i,r-1}\leq\Delta_{th})$, which means device $i$ cannot make good enough contributions to FL training with current size of subnetwork, and should adaptively increase the subnetwork size from training efficiency perspective. Thus, following Eqn. (4.1), $TE_{i,r}=TE_{i,r-1}+\gamma$; Case II: $(l_{i,r-1}\geq\alpha\cdot l_{target})$ and $(\Delta l_{i,r-1}>\Delta_{th})$, which means after increasing the subnetwork size in Case I, device $i$ can make good contributions to FL training again, and should keep the subnetwork size from training efficiency perspective. Thus, following Eqn. (4.1), $TE_{i,r}=TE_{i,r-1}$. Device $i$ may switch between those two cases at the middle FL training stage. At the late FL training stage when FL is about to converge, $(l_{i,r-1}<\alpha\cdot l_{target})$, i.e., $l_{i,r-1}$ stabilizes at a small value, and $(\Delta l_{i,r-1}\leq\Delta_{th})$. That indicates device $i$ has already made substantial contributions to FL training and sticking to the large size subnetwork for $i$’s local training has very limited benefits for FL convergence. Thus, device $i$ can gradually downsize its local subnetworks from training efficiency perspective, i.e., $TE_{i,r}=TE_{i,r-1}-\gamma$. As such, Eqn. (4.1) can characterize FL training dynamics and adjust the training efficiency utility values for devices’ adaptive subnetwork selection at different training stages.
$\begin{split}Util(i,r)=\underbrace{\left({T_{i,r}^{tr}+T_{i,r}^{co}}\right)^{-1}}_{\textit{System efficiency utility}}\times\,\,\,{\underbrace{\left\{TE_{i,r-1}+\gamma\cdot\left[2\cdot\mathbbm{1}(l_{i,r-1}\geq\alpha\cdot l_{target})-1\right]\cdot\mathbbm{1}(\Delta l_{i,r-1}\leq\Delta_{th})\right\}}_{\textit{Training efficiency utility}}}^{\beta}\end{split}$ (5)
Adaptive subnetwork selection utility function. WHALE-FL trades-off the system and training efficiencies to determine the utility values for subnetwork scheduling over rounds. The adaptive subnetwork selection utility function is shown in Eqn. (5), where $Util(i,r)$ associates system and training efficiencies with developer-specified factor $\beta$. Aware of both system and FL training dynamics, a large/small value of $Util(i,r)$ suggests that device $i$ should opt for a large/small sized subnetwork in the subsequent $r$-th round.
4.2 Utility Value to Subnetwork Size Conversion
The calculated utility in Eqn. (5) cannot directly be used by individual mobile device to decide its subnetwork size selection. To facilitate mobile devices’ decisions, it is necessary to convert subnetwork selection utility values into available/candidate subnetwork sizes.
Given definitions above, the next step is to normalize devices’ utility values into the range of $[0,1]$, in order to identify the model shrinkage ratio. We propose to use a piecewise linear function to normalize $Util(i,r)$ into $U_{n}(i,r)$ as follows.
$U_{n}(i,r)=\begin{cases}\frac{Util(i,r)}{U_{th}},&Util(i,r)\leq U_{th},\\ 1,&\text{otherwise},\end{cases}$ (6)
where $U_{th}$ is a configurable threshold that represents the utility level at which the full-sized model should be adopted.
After the utility value normalization, device $i$ selects its subnetwork for the $r$-th round local training by
$W(i,r)=\begin{cases}\hat{W}(i,r),&\textbf{if }|W^{max}_{i}|>|\hat{W}(i,r)|;\\ W^{max}_{i},&\textbf{if }|W^{max}_{i}|\leq|\hat{W}(i,r)|.\end{cases}$ (7)
Here, $|W^{max}_{i}|$ denotes the maximum subnetwork size that device $i$ can support with its full computing capacity, where $W^{max}_{i}\in\mathcal{W}^{P}$ as defined in Sec. 2.2. $\hat{W}(i,r)\in\mathcal{W}^{P}$ is a subnetwork derived from normalized utility value $U_{n}(i,r)$, which can be expressed as
$\hat{W}(i,r)=\begin{cases}W^{1},&\textbf{if }U_{n}(i,r)\geq\frac{(P-1)}{P};\\ W^{2},&\textbf{if }U_{n}(i,r)\in[\frac{(P-2)}{P},\frac{(P-1)}{P});\\ \cdots,&\cdots\\ W^{p},&\textbf{if }U_{n}(i,r)\in[\frac{(P-p)}{P},\frac{(P-p+1)}{P});\\ \cdots,&\cdots\\ W^{P},&\textbf{if }U_{n}(i,r)<\frac{1}{P},\end{cases}$ (8)
where $|W^{p}|/|W_{g}|=s^{2(p-1)},\forall W^{p}\in\mathcal{W}^{P}$.
Algorithm 1 WHALE-FL Procedure
1:Input: Data $\{D_{1},\cdots,D_{M}\}$ distributed on $M$ mobile devices, the number of local epochs $E$, the local minibatch size $B$, the learning rate $\eta$, the channel shrinkage ratio $s$, and the number of subnetwork size levels $P$.
2:Initialize global model $W(g,0)$
3:for each communication round $r=0,1,2,\cdots$ do
4:     for each client $i\in[M]$ in parallel do
5:         $s_{i}\leftarrow$ SubnetSelect$(SE_{i,r},TE_{i,r},W_{i}^{max},P)$
6:         $d_{i}\leftarrow s_{i}d_{g},k_{i}\leftarrow s_{i}k_{g}$
7:         $W(i,r)\leftarrow W(g,r)[:d_{i},:k_{i}]$
8:         $W(i,r+1)\leftarrow$ ClientUpdate$(i,s_{i},W(i,r))$
9:     end for
10:     for each size level $p\in[P]$ do
11:         Compute $W_{g}^{p,r+1}$ via Eqn. (2.2)
12:     end for
13:     Update the global model $W(g,r+1)$ via Eqn. (9)
14:end for
15:SubnetSelect$(SE_{i,r},TE_{i,r},W_{i}^{max},P)$:
16:if $r=0$ then
17:     $s_{i}\leftarrow 1$
18:else
19:     $Util(i,r)\leftarrow SE_{i,r}\times\left(TE_{i,r}\right)^{\beta}$
20:     Compute $U_{n}(i,r)$ via Eqn. (6)
21:     Compute $\hat{W}(i,r)$ via Eqn. (8)
22:     Compute $W(i,r)$ via Eqn. (7) and determine the corresponding $p$
23:     $s_{i}\leftarrow s^{p-1}$
24:end if
25:Return $s_{i}$
26:ClientUpdate$(i,s_{i},W(i,r))$:
27:$B_{i}\leftarrow$ Split local data $D_{i}$ into batches of size $B$
28:for each local epoch $e$ from $1$ to $E$ do
29:     for batch $b_{i}\in B_{i}$ do
30:         $W(i,r+1)\leftarrow W(i,r)-\eta\nabla l(W(i,r),s_{i};b_{i})$
31:     end for
32:end for
33:Return $W(i,r+1)$
Then, mobile devices conduct local computing according to their selected subnetworks, respectively, followed by transmitting local model updates to FL server. Following the same aggregation method in [9], FL server aggregates updated local models with heterogeneous subnetworks and updates the global model as
$W(g,r+1)=W_{g}^{P,r+1}\cup\bigcup^{P}_{p=2}W_{g}^{p-1,r+1}\setminus W_{g}^{p,r+1}.$ (9)
In summary, during FL training, mobile devices collect their local information at runtime, including up-link channel quality, background computational loads, memory usage, training loss, etc. Based on the collected information, at the beginning of the $r$-th training round, each device leverages Eqn. (5) to trade-off system efficiency and training efficiency, and calculates its adaptive subnetwork selection utility value $Util(i,r)$. The utility value is then normalized into $U_{n}(i,r)$. Device $i$ uses $U_{n}(i,r)$ to determine the subnetwork size and select an appropriate subnetwork for its local training according to Eqn. (7) and Eqn. (8). After that, FL server aggregates locally trained subnetworks with different sizes and updates the global model for the next round training. The pseudocode of WHALE-FL procedure is provided in Alg. 1.
5 Experimental Setup
5.1 WHALE-FL Testbed
The WHALE-FL testbed consists of an FL aggregator and a set of heterogeneous mobile devices as FL clients. A NVIDIA RTX 3090 serves as the FL server, whose memory capacity is $24$ GB. For heterogeneous FL clients, we have incorporated $5$ types of mobile devices, i.e., MacBookPro2018, NVIDIA Jetson Xavier, NVIDIA Jetson TX2, NVDIA Jetson Nano, and Raspberry Pi 4, representing a range of on-device computing capabilities from high to low. The WHALE-FL system involves a total of $20$ mobile devices, $4$ devices per type. Communication between FL clients and the FL server is facilitated through LTE, BlueTooth, and Wi-Fi $5$ transmission environments. The corresponding transmission rates are $80$ Mbps (Wi-Fi $5$), $20$ Mbps (LTE), and $10$ Mbps (BlueTooth $3.0$), respectively. We set hidden channel shrinkage ratio $s=\frac{1}{2}$ and adopt $5$ subnetwork size levels. Accordingly, the model shrinkage ratios for the $5$ size levels (i.e., $p=1,2,\cdots,5$) are $1$, $\frac{1}{4}$, $\frac{1}{16}$, $\frac{1}{64}$, and $\frac{1}{256}$, respectively.
5.2 Datasets, Models, Parameters and Baselines
We conduct our experiments with three different FL tasks: image classification, human activity recognition and language modeling. As for the image classification task, we train a CNN on MNIST dataset [7] and a ResNet18181818 on CIFAR10101010 dataset [12]. Human activity recognition involves training a CNN on the HAR dataset [10], and a Transformer is trained on the WikiText2 dataset [8] for the language modeling task. We use the balanced non-IID data partition [16]. Take the MNIST dataset as example, the total number of classes is $10$. Our default setup is that each device has $\sigma=2$ classes. We apply similar non-IID setup to other tasks. Besides, we set $\Delta_{th}$ to follow the typical learning rate. As for $l_{target}$, it is the only task-specific hyperparameter. We let $l_{target}$ be $0.001$ for CNN@MNIST and CNN@HAR, $0.1$ for Resnet18181818@CIFAR10101010 and $1$ for Transformer@WikiText2222 in our experiments.
We compare our WHALE-FL against two peer designs across these FL tasks: i) FedAvg [20], where all the clients train with full-sized models, and ii) HeteroFL [9], where subnetwork assignments are fixed and align with clients’ full computation and communication capabilities.
Figure 2: Performance comparison of different FL training approaches under various learning tasks. Figures from left to right are CNN@MNIST, ResNet18181818@CIFAR10101010, Transformer@WikiText2222, and CNN@HAR with non-IID datasets.
6 Evaluation and Analysis
6.1 Latency Efficiency and Learning Performance
As the results shown in Fig. 2, the proposed WHALE-FL consistently achieves remarkable training speedup across various FL tasks without sacrificing learning accuracy. Compared with FedAvg, WHALE-FL accelerates the FL training to the target testing accuracy by approximately 1.5x, 1.9x, 1.3x and 2.1x for FL tasks including CNN@MNIST, ResNet18181818@CIFAR10101010, Transformer@WikiText2, and CNN@HAR, respectively. As detailed in Sec. 3, HeteroFL’s static fixed-size subnetwork assignment policy is not aware of system and training dynamics, which may slow down FL convergence. In contrast, considering both system efficiency and training efficiency, WHALE-FL appropriately assesses the subnetwork selection utility for individual device and adaptively adjusts the local subnetwork size to suit for time-varying communication and computational conditions and dynamic changing requirements of FL training at different FL training stages, in order to reduce training latency. Consequently, compared with HeteroFL, WHALE-FL achieves a notable speedup of 1.2x, 1.3x, 1.2x and 1.5x for the tested 4 learning tasks, respectively.
Figure 3: System efficiency utility and subnetwork size level changes over training time (Macbookpro2018, CNN@HAR).
Figure 4: Performance comparison: WHALE-FL, system efficiency only and training efficiency only designs (ResNet18181818@CIFAR10101010).
6.2 Subnetwork Size Changes over Rounds
As shown in Fig. 3, the subnetwork size for Macbookpro2018 (CNN@HAR) dynamically changes during FL training process. At the early training stage, the subnetwork size changes along with system efficiency utility. At the middle training stage, the subnetwork size gradually increases jointly determined by system and training efficiency. At the late stage, the subnetwork size reduces for better training latency efficiency. Here, we take Macbookpro2018 for example, and the analysis applies to all participating mobile devices.
6.3 System Efficiency vs Training Efficiency
To differentiate system efficiency’s contributions from training efficiency’s ones, we compare WHALE-FL with system efficiency utility only and training efficiency utility only schedulings. As the results shown in Fig. 4, WHALE-FL converges faster than training efficiency only subnetwork scheduling when achieving the target accuracy, since training efficiency only design has no consideration of system dynamics and its impacts on subnetwork size selection; WHALE-FL has better testing accuracy but proceeds slower than system efficiency only subnetwork scheduling at the early training stage. The reason behind is that system efficiency only design prioritizes system dynamics while ignoring dynamic model accuracy requirements for local training at different FL training stages. WHALE-FL trades-off system and training efficiencies and jointly considers their benefits for FL training.
Figure 5: Sensitivity performance under different $\alpha$, $\beta$, and $\gamma$ value configurations (CNN@MNIST).
Figure 6: Sensitivity performance under different $\alpha$, $\beta$, and $\gamma$ value configurations (Transformer@WikiText2).
6.4 Sensitivity Analysis
We further evaluate the impacts of $\alpha$, $\beta$ and $\gamma$, defined in the subnetwork selection utility function, on subnetwork scheduling. Taking CNN@MNIST as an example, we conduct sensitivity analysis of its performance under different $\alpha$, $\beta$ and $\gamma$ values, and present the results in Fig. 5. For generalization purpose, we have also conducted the sensitivity study for NLP task, whose results are shown in Fig. 6.
Figure 7: Sensitivity performance under different $U_{th}$ value configurations (Transformer@WikiText2222).
The $\alpha$-value helps to separate the middle and late FL training stages, which determines the increment/decrement of selected subnetwork sizes. If $\alpha=1$, FL training stops when ($l_{i,r-1}\leq l_{target}$) as shown in Eqn. (5) and there will be no late stage and corresponding subnetwork size decrease. Thus, as shown in Fig. 5, compared with $\alpha=2$ or $5$ cases, $\alpha=1$ based FL training is a little bit slow to converge.
The hyperparameter $\beta$ trades-off system efficiency and training efficiency utilities. The large/small $\beta$ value means that the device prioritizes training/system efficiency. As the results shown in Fig. 5, we find that the FL training converges slower but achieves higher testing accuracy when $\beta$ is large, e.g., $\beta=5$ in Fig. 5, while FL training is faster at early stage but achieves lower testing accuracy when $\beta$ is smaller, e.g., $\beta=1$ in Fig. 5. System efficiency and training efficiency are somehow balanced when $\beta=2$ in Fig. 5. Thus, although $\beta$ is a developer-specified factor, a proper selection of $\beta$ value helps FL training converge fast while achieving good learning performance.
Figure 5 reveals the training performance is not that sensitive to $\gamma$ values. Recall that $\gamma$ is the step-size for increment/decrement of the training efficiency utility. The results reflect that our WHALE-FL consistently outperforms the baseline, FedAvg, regardless of how the increment/decrement step-size is configured in Eqn. (5). This may be attributed to the proposed normalization and discretization operations on utility values in Sec. 4.2, which mitigates the impacts of $\gamma$ on the latency or learning performance. For developer-specified $U_{th}$, it can serve as a cross-device guidance parameter for selecting subnetwork sizes. The impact of $U_{th}$ is shown in Fig. 7. Increasing $U_{th}$ makes all clients opt for smaller subnetworks, and decreasing $U_{th}$ makes all clients tend to select bigger subnetworks.
6.5 Impacts of Data Heterogeneity
We further evaluate the impacts of data heterogeneity on WHALE-FL’s performance. Here, we take CNN@MNIST as an example and use the balanced non-IID data partition [16]. The total number of classes in the MNIST dataset is $10$. We study the cases that each device has $\sigma=2,5$ or $10$ classes, where the data distribution is IID if $\sigma=10$, i.e., every device has all classes. The results are shown in Table 1, where we find that (i) FL training with non-IID data takes longer time to converge, and (ii) embracing both system and training efficiency utilities, WHALE-FL can remarkably improve FL training delay efficiency compared with HeteroFL and FedAvg under various data heterogeneity scenarios.
Table 1: Performance comparison under different data heterogeneity levels (CNN@MNIST), where “OL" represents the overall latency.
Local Model CNN@MNIST
non-IID Level $\sigma=2$ $\sigma=5$ $\sigma=10$
Target Acc. 90% 95% 97%
Metric Hours (OL) Hours (OL) Hours (OL)
WHALE 0.82 (1.5x) 0.24 (1.4x) 0.08 (3.8x)
HeteroFL 0.98 (1.3x) 0.26 (1.3x) 0.10 (3.0x)
FedAvg 1.23 (1.0x) 0.33 (1.0x) 0.30 (1.0x)
7 Conclusion
In this paper, we have proposed WHALE-FL, a wireless and heterogeneity aware latency efficient federated learning approach, to accelerate FL training over mobile devices via subnetwork scheduling. Unlike existing static fixed-size subnetwork assignments, WHALE-FL has incorporated an adaptive subnetwork scheduling policy, enabling mobile devices to flexibly select subnetwork sizes for local training, with a keen awareness of mobile devices’ system dynamics and FL training dynamics. At its core, WHALE-FL has employed a well-designed subnetwork selection utility function, capturing changes in the device’s system conditions (including available computing and communication capacities) and evolving FL training requirements for local training, to schedule appropriate subnetworks for mobile devices in each FL training round. Results from prototype-based evaluations have demonstrated that WHALE-FL surpasses peer designs, significantly accelerating FL training over heterogeneous mobile devices without sacrificing learning accuracy.